OpenStudy (anonymous):

Solve using substitution: 2x + y = 3 -3y =12 − x Answer (3, -3) (4, -1) No solution Infinitely many solutions

OpenStudy (anonymous):

I prefer not to use substitution. If you have to though, you can take the given points and put them in for x and y in your equations.

OpenStudy (istim):

Rearrange one of the eq'ns, and then 'substitute' it for one of the variables of the other eq'n.

OpenStudy (istim):

If you need a more detailed explanation, or need to double check your work, call out.

OpenStudy (anonymous):

ii dnt kn hw to do it

OpenStudy (anonymous):

Take the given points (3, -3). 3 is your x and -3 is your y. Plug that in to 2x + y = 3. 2(3)+(-3)=3 6-3=3 3=3

OpenStudy (anonymous):

what if this was a problem Solve using elimination: -5x = 3y − 7 10x + 6y = 14 Answer (-4, 9) (4, -1) Infinitely many solutions No solution

OpenStudy (anonymous):

For elimination, you need to either get rid of the x or the y